It takes the Canon printer 15 min longer to print a set of photos than it takes the HP printer. Together, it takes them 180/7 min to print the photos. How long would it take each machine, working along, to print the photos?
Answer by richwmiller(17219)

1/x+1/(x+15)=7/180
180/x+180/(x+15)=7
x=45 x+15=60